A man drove 15/4 miles from his home to his office . he went out for a lunch for a near by restaurant 2/3 miles from his office and back . later on he goes to a store 2 miles from his office and then drove to home 11/6 . what distance in total did he travel ?

Answer by jorel1380(3719)

15/4 + 2(2/3)+ 2 + 11/6=
45/12+16/12+24/12+22/12=????
☺☺☺☺